OpenStudy (anonymous):

Please Help. Will Give Medal!!! Suppose you had 0.70 g of Ba(NO3)2 and 1.24 g of NH2SO3H, which would be the limiting reactant in reactions (a), (b) and (c)? Is the limiting reactant the same for all three? a. _____ Ba(NO3)2(aq) + _2_ NH2SO3H(aq) -> _____ Ba(NH2SO3)2(s) + _2_ HNO3(aq) b. _____ Ba(NO3)2(aq) + _____ NH2SO3H(aq) + ____ H2O(l) -> ____ BaSO4(s) + ____ NH4NO3(aq) + ___ HNO3(aq) c. ____ Ba(NO3)2(aq) + _2_ NH2SO3H(aq) + _2_ H2O(l) -> ____ Ba(NH2)2(s) + _2_ H2SO4(aq) + _2_ HNO3(aq)

OpenStudy (anonymous):

Only the question is there...

OpenStudy (cuanchi):

you have to calculate how many moles of each of the reactants you have in the given grams, then look at the relationship of the reactants in the 3 reactions, the a) is 1:2, b) 1:1 and c)is 1:2.

OpenStudy (anonymous):

i dont understand how to convert to moles

OpenStudy (cuanchi):

number of moles = mass/ molecular mass you have to calculate the molecular mass of your compounds do you know how to calculate molecular mass?

OpenStudy (anonymous):

no

OpenStudy (cuanchi):

do you have a periodic table in your hands?

OpenStudy (anonymous):

yes

OpenStudy (cuanchi):

find there the atomic mass of all the elements that compose your first compound, Ba, N, O Ba(NO3)2

OpenStudy (anonymous):

okay. then what do i do with those numbers?

OpenStudy (cuanchi):

the atomic mass is the 4-5 digit number in the box of the element for Ba = 137.327, N= 14.0067 , O= 15.9994

OpenStudy (cuanchi):

you are good in math! you have to add this number as the proportions in the formula

OpenStudy (cuanchi):

1 Ba + 2 x (1N +3 x O) =

OpenStudy (anonymous):

im confused

OpenStudy (cuanchi):

137.327+ 2 (14.0067 +( 3 x 15.9994))= 261.337

OpenStudy (anonymous):

what is the 261.337? i'm sorry, i just don't understand this problem.

OpenStudy (cuanchi):

that is the molecular mass of Ba(NO3)2

OpenStudy (cuanchi):

it is equal to the mass of one mol of Ba(NO3)2 in grams

OpenStudy (anonymous):

okay. so i find the molar mass of Ba(NO3)2 and NH2SO3H, then what do i do?

OpenStudy (cuanchi):

then calculate how many mol of each compound do you have in the 0.70 g of Ba(NO3)2 and 1.24 g of NH2SO3H

OpenStudy (anonymous):

how do i do that?

OpenStudy (cuanchi):

you have to divide the "mass/molecular mass" 0.70/261.337=0.002678

OpenStudy (anonymous):

okay

OpenStudy (cuanchi):

for the reaction a) you need one mole of Ba(NO3)2 every 2 mol of NH2SO3H 1:2 relationship

OpenStudy (anonymous):

okay. and is Ba(NO3)2 the limiting reactant in all 3?

OpenStudy (cuanchi):

and you have a 1:5 relationship 0.002678/0.01277, that means the you have excess of NH2SO3H and your limiting reactant is Ba(NO3)2

OpenStudy (cuanchi):

I think so...

OpenStudy (anonymous):

okay thank you for helping me!